# Cross-Modal Safety Mechanism Transfer in Large Vision-Language Models

### v1

@@ Abstract
Vision-language alignment in Large Vision-Language Models (LVLMs) successfully enables LLMs to understand visual input. However, we find that existing vision-language alignment methods fail to transfer the existing safety mechanism for text in LLMs to vision, which leads to vulnerabilities in toxic image. To explore the cause of this problem, we give the insightful explanation of where and how the safety mechanism of LVLMs operates and conduct comparative analysis between text and vision. We find that the hidden states at the specific transformer layers play a crucial role in the successful activation of safety mechanism, while the vision-language alignment at hidden states level in current methods is insufficient. This results in a semantic shift for input images compared to text in hidden states, therefore misleads the safety mechanism. To address this, we propose a novel Text-Guided vision-language Alignment method (TGA) for LVLMs. TGA retrieves the texts related to input vision and uses them to guide the projection of vision into the hidden states space in LLMs. Experiments show that TGA not only successfully transfers the safety mechanism for text in basic LLMs to vision in vision-language alignment for LVLMs without any safety fine-tuning on the visual modality but also maintains the general performance on various vision tasks (Safe and Good).
@@ 1 Introduction
Vision-language alignment methods for Large Vision-Language Models (LVLMs) use a basic LLM, a lightweight vision encoder and projector to efficiently enable the LLM to understand visual input for various vision tasks with relatively low training costs (Liu et al., 2024c; Dai et al., 2023; Zhu et al., 2023). Recent studies indicate the safety of LVLMs deserves attention (Liu et al., 2024a; Wang et al., 2023; Gong et al., 2023). Given that vision and language are aligned into a common space in LVLMs, the safety mechanism should be shared by both of them. However, this is not the case. We find that compared to toxic text input, LVLMs are more vulnerable to toxic vision input. Existing studies improve the safety of LVLMs by utilizing toxic vision data for safety fine-tuning (Zong et al., 2024; Wang et al., 2024). However, they do not provide an essential explanation for the question: “Why can’t the safety mechanism for text be shared by vision after vision-language alignment?”.
@@
Since the mainstream vision-language alignment methods for LVLMs (e.g., LLaVA (Liu et al., 2024c)) lack additional safety fine-tuning and the training data in this process contain few toxic samples (Wang et al., 2023), the safety mechanism of LVLMs is mostly inherited from the safety mechanism that have established for text in their basic LLMs. So an intuitive view to explain the dilemma of LVLMs’ safety on vision is that the vision-language alignment cannot effectively transfer the safety mechanism for text in LLMs to vision. Based on this, this paper proposes a novel perspective called Cross-Modal Safety Mechanism Transfer to rethink, explain and address the exacerbated vulnerability of LVLMs to toxic vision input compared to toxic text input. Cross-modal safety mechanism transfer means transferring the existing safety mechanism for text in LLMs to vision in vision-language alignment training without any additional safety fine-tuning on vision. Our experiments in § 3 reveal that current vision-language alignment methods fail to achieve an effective cross-modal safety mechanism transfer. LVLMs exhibit significantly different safety capabilities when handling toxicity with the same semantics but different modalities (as shown in Figure 1).
@@
To analyze the cause of this issue, which also provides an essential explanation for the vulnerability of LVLMs to toxic vision input, we first offer an insightful understanding of where and how the safety mechanism of LVLMs operates in § 4.1. Specifically, we propose a novel method to locate the activation of safety mechanism in LVLMs and analysis the attention pattern on toxic tokens. We find that the hidden states at the specific local transformer layers play a crucial role in the successful activation of safety mechanism (Figure 1a). Then our comparative analysis between text and vision modalities in § 4.2 reveals that current vision-language alignment methods in training LVLMs cannot effectively align the vision with the language at hidden states level (Figure 1b). This misalignment causes that the hidden states of the vision shifts from the hidden states of the corresponding text with the same semantics. When the hidden states of toxic image are input to the specific transformer layers responsible for safety mechanism activation, this shift prevents these layers from correctly capturing the semantics of image as they do for text, thereby hindering their ability to accurately assess the toxicity in images (Figure 1c). This is the key reason why the safety mechanism for text in basic LLMs cannot be transferred to vision in vision-language alignment training for LVLMs.
@@
To address the above problem, we propose a novel vision-language alignment training method called TGA, which uses retrieved texts related to the input vision to guide the projection of vision into the hidden states space in LLMs, thereby achieving alignment at hidden states level. Extensive experiments show that TGA successfully transfers the safety mechanism for text in basic LLMs to vision during vision-language alignment training for LVLMs without any safety fine-tuning on the visual modality. Besides, our TGA maintains general performance across various vision tasks compared with existing state-of-the-art (SoTA) LVLMs, such as InstructBLip, LLaVA-1.5, Qwen-VL-Chat in VQA, SEED-Bench, MM-Vet, etc. This indicates our TGA is a safe and effective vision-language alignment method for training LVLMs. The contributions of this paper are as follows:
@@
  • We propose a novel perspective called Cross-Modal Safety Mechanism Transfer, which aims to transfer the safety mechanism for text in LLMs to vision in vision-language alignment without any safety fine-tuning on the vision data. This rethinks, explains, and addresses the exacerbated vulnerability of LVLMs to toxic vision compared to toxic text.
@@
  • We explain where and how safety mechanism in LVLMs operates, and reveal that current vision-language alignment methods fail to achieve effective cross-modal safety mechanism transfer because of insufficient alignment at hidden states level.
@@
  • We propose a novel vision-alignment method called TGA that can not only transfer the safety mechanism against toxic text in basic LLMs to vision but also maintain the general performance on various vision tasks compared with existing SoTA LVLMs.
@@ 2 Related Work
Vision-Language Alignment in LVLMs. Vision-language alignment in LVLMs equips basic LLMs with the ability to understand and process visual input by pre-training and instruction-tuning on large-scale text-image pairs such as works in LLaVA (Liu et al., 2024c), InstructBLip (Dai et al., 2023), Qwen-VL (Bai et al., 2023b), MiniGPT-4 (Zhu et al., 2023), Flamingo (Awadalla et al., 2023), PaLM-E (Driess et al., 2023), etc. However, it remains unknown whether visual-language alignment can extend all the capabilities of LLM on text to vision. This paper unveils that the safety mechanism of LLMs on text cannot be transferred to vision in existing typical vision-language alignment methods. We give the essential explanation for the cause of this issue and the novel solution method.
@@ 3 Safety Mechanism Cannot be Cross-Modal Transferred
In this part, we provide LVLM and its basic LLM with toxic inputs having the same semantics but in different modalities (text and vision) to evaluate the safety capabilities on different modalities, which is used as an assessment of cross-modal safety mechanism transfer. We find that safety mechanism for text is not effectively transferred to vision in visual-language alignment.
Data Construction. We collect real toxic images from open-source datasets. For each image, we use LLaVA-NEXT (Liu et al., 2024b) to generate caption for it to get the toxic text-image pair. Text and image in this pair have the same semantics but are in different modalities. The specific datasets include HOD (Ha et al., 2023) that contains $10,631$ toxic images about alcohol, cigarette, gun, insulting gesture and knife, and ToViLaG (Wang et al., 2023) that contains $9,900$ toxic images about bloody and porn. After the caption generation, we get 20,31 toxic text-image pairs for experiments.
@@
Findings. The experimental results are shown in Table 1. Different LVLMs across different toxic scenes show the following common conclusions: (1) DSR of LVLM and its basic LLM is close on toxic text, it indicates that safety mechanism of basic LLMs on text is successfully preserved in visual-language alignment training of LVLMs. (2) For the toxic information with the same semantics in different modalities (text and vision), the safety capabilities of LVLMs vary significantly, and LVLMs can hardly defense toxicity in the visual modality. It indicates that safety mechanism for text is not effectively transferred to vision in visual-language alignment training of LVLMs.
@@ 4 Cause of Failure in Cross-Modal Transferring Safety
This section analyzes the cause of the failure in transferring the safety mechanism from text to vision. Firstly, in § 4.1, we find that hidden states at the specific transformer layers play a crucial role in the successful activation of safety mechanism. Then, our comparative analysis between text and image in § 4.2 reveals that alignment between vision and language at hidden states level is insufficient. It makes the transformer layers responsible for safety mechanism activation cannot correctly capture the semantics of image as they perform on text, so they cannot correctly assess the toxicity in image and the safety mechanism on vision collapses.
@@ 4.1 Safety Mechanism is Activated at Specific Layers by Hidden States
This part gives an insightful explanation of where and how the safety mechanism activated in LVLMs on textual, which is fundamental to understand the safety mechanism collapse on vision. Previous studies (Tenney, 2019; Dai et al., 2021; Meng et al., 2022) have shown that transformers in different layers of language model have different functions such as lexical, semantic, knowledge, etc., this paper proposes a novel method to identify the transformer layers responsible for activating safety mechanism in LVLMs and analyzes the attention patterns on toxic tokens within these layers to give the explanation of where and how the safety mechanism is activated.
@@
in which $\textrm{{W}}\in\mathbb{R}^{h\times v}$ is the vocabulary head that maps the output hidden states $\textrm{{H}}^{\prime}$ to the word distribution in vocabulary $\mathcal{X}$ with size $v$. Since previous studies (Chuang et al., 2024; Xu et al., 2024) prove that due to the residual connections, the combination of any intermediate layer in a language model with the vocabulary head W can represent its distribution of semantics, we apply W to each intermediate layer to obtain its distribution of semantics as:
@@
Sorry tokens start to show the most significant growth among the whole vocabulary means that the transformer layer realizes that the current input is toxic, so it tries to update the word distribution to refuse following the instruction, which is actually the safety mechanism activation. So our method makes sense. Safety mechanism perturbation experiments in Figure 3 also show that our method can accurately determine the activation location of the safety mechanism of LVLMs.
@@ 5.1 Motivation and Overview
As shown in Figure 6 (a), previous vision alignment methods in LVLMs construct image-instruction-output triples as training data, and use cross-entropy in language modeling as the loss fuction to optimize the output generated by basic LLMs to make LLM understand the input vision, thereby achieving visual-language alignment. We name these methods input-to-output alignment, i.e. align the text output to the visual input, which is actually asymmetrical. The natural flaw of these method is that they black-box basic LLMs and only focus on output, while neglecting whether the internal representation of the visual input in LLMs, i.e., the hidden states, aligns with the hidden states in text modality. To address this problem, we propose the Text-Guided input-to-input Alignment (TGA) at hidden states level, as shown in Figure 6 (b). For the input image, TGA retrieves the semantically relevant text as the template to guide the alignment of vision to language at hidden states level.
@@ A.1 General Capabilities of LVLMs on Vision Tasks
(1) ScienceQA (Lu et al., 2022): A benchmark that consists of  21k multimodal multiple choice questions with a diverse set of science topics and annotations of their answers with corresponding lectures and explanations. We follow LLaVA (Liu et al., 2024c) to evaluate the zero-shot generalization of LVLMS on scientific question answering in image subset and use accuracy as the metric.
@@
A.2 Case Study

### v2

@@ Abstract
Content warning: This paper contains harmful images and texts!
Vision-language alignment in Large Vision-Language Models (LVLMs) successfully enables LLMs to understand visual input. However, we find that existing vision-language alignment methods fail to transfer the existing safety mechanism for text in LLMs to vision, which leads to vulnerabilities in toxic image. To explore the cause of this problem, we give the insightful explanation of where and how the safety mechanism of LVLMs operates and conduct comparative analysis between text and vision. We find that the hidden states at the specific transformer layers play a crucial role in the successful activation of safety mechanism, while the vision-language alignment at hidden states level in current methods is insufficient. This results in a semantic shift for input images compared to text in hidden states, therefore misleads the safety mechanism. To address this, we propose a novel Text-Guided vision-language Alignment method (TGA) for LVLMs. TGA retrieves the texts related to input vision and uses them to guide the projection of vision into the hidden states space in LLMs. Experiments show that TGA not only successfully transfers the safety mechanism for text in basic LLMs to vision in vision-language alignment for LVLMs without any safety fine-tuning on the visual modality but also maintains the general performance on various vision tasks. Code is available [^1] https://github.com/xsc1234/VLM_Safety_Transfer.
@@ 1 Introduction
Vision-language alignment methods for Large Vision-Language Models (LVLMs) use a basic LLM, a lightweight vision encoder and projector to efficiently enable the LLM to understand visual input for various vision tasks with relatively low training costs (Liu et al., 2024c; Dai et al., 2023; Zhu et al., 2023). Recent studies indicate that the safety of LVLMs deserves attention (Liu et al., 2024a; Wang et al., 2023; Gong et al., 2023). Given that vision and language are aligned into a common space in LVLMs, the safety mechanism should be shared by both of them. However, this is not the case. We find that compared to toxic text input, LVLMs are more vulnerable to toxic vision input. Existing studies on the safety of LVLMs (Zong et al., 2024; Wang et al., 2024) fall short of providing an essential explanation for the question: “Why can’t the safety mechanism for text be shared by vision after vision-language alignment?”. This is the core issue that this paper aims to address.
@@
Since the mainstream vision-language alignment methods for LVLMs (e.g., LLaVA (Liu et al., 2024c)) lack additional safety fine-tuning and the training data in this process contain few toxic samples (Wang et al., 2023), the safety mechanism of LVLMs is mostly inherited from the safety mechanism that has been established for text in their basic LLMs. So an intuitive view to explain the dilemma of LVLMs’ safety on vision is that the vision-language alignment cannot effectively transfer the safety mechanism for text in LLMs to vision. Based on this, this paper proposes a novel perspective called Cross-Modal Safety Mechanism Transfer to rethink, explain and address the exacerbated vulnerability of LVLMs to toxic vision inputs compared to toxic text inputs. Cross-modal safety mechanism transfer means transferring the existing safety mechanism for text in LLMs to vision in vision-language alignment training without any additional safety fine-tuning on vision. Our experiments in § 3 reveal that current vision-language alignment methods fail to achieve an effective cross-modal safety mechanism transfer. LVLMs exhibit significantly different safety capabilities when handling toxicity with the same semantics but different modalities (as shown in Figure 1).
@@
To analyze the cause of this issue, which also provides an essential explanation for the vulnerability of LVLMs to toxic vision input, we first offer an insightful understanding of where and how the safety mechanism of LVLMs operates in § 4.1. Specifically, we propose a novel method to locate the activation of safety mechanism in LVLMs and analyze the attention pattern on toxic tokens. We find that the hidden states at the specific local transformer layers play a crucial role in the successful activation of safety mechanism (Figure 1a). Then our comparative analysis between text and vision modalities in § 4.2 reveals that current vision-language alignment methods in training LVLMs cannot effectively align the vision with the language at hidden states level (Figure 1b). This misalignment causes the hidden states of the vision to shift from the hidden states of the corresponding text with the same semantics. When the hidden states of toxic image are input to the specific transformer layers responsible for safety mechanism activation, this shift prevents these layers from correctly capturing the semantics of images as they do for texts, thereby hindering their ability to accurately assess the toxicity in images (Figure 1c). This is the key reason why the safety mechanism for text in basic LLMs cannot be transferred to vision in vision-language alignment training for LVLMs.
@@
To address the above problem, we propose a novel vision-language alignment training method called TGA, which uses retrieved texts related to the input vision to guide the projection of vision into the hidden states space in LLMs, thereby achieving alignment at the hidden-state level. Extensive experiments show that TGA successfully transfers the safety mechanism for text in basic LLMs to vision during vision-language alignment training for LVLMs without any safety fine-tuning on the visual modality. Besides, our TGA maintains general performance across various vision tasks compared with existing state-of-the-art (SoTA) LVLMs. This indicates our TGA is a safe and effective vision-language alignment method for training LVLMs. The contributions of this paper are:
@@
$\bullet$ We propose a novel perspective called Cross-Modal Safety Mechanism Transfer, which aims to transfer the safety mechanism for text in LLMs to vision in vision-language alignment without any safety fine-tuning on the vision data. This rethinks, explains, and addresses the exacerbated vulnerability of LVLMs to toxic vision compared to toxic text.
@@
$\bullet$ We explain where and how safety mechanism in LVLMs operates, and reveal that current vision-language alignment methods fail to achieve effective cross-modal safety mechanism transfer because of insufficient alignment at hidden states level.
@@
$\bullet$ We propose a novel vision-alignment method called TGA that can not only transfer the safety mechanism against toxic text in basic LLMs to vision but also maintain the general performance on various vision tasks compared with existing SoTA LVLMs.
@@ 2 Related Work
Vision-Language Alignment in LVLMs. Vision-language alignment in LVLMs equips basic LLMs with the ability to understand and process visual input by pre-training and instruction-tuning on large-scale text-image pairs such as works in LLaVA (Liu et al., 2024c), InstructBLip (Dai et al., 2023), Qwen-VL (Bai et al., 2023b), MiniGPT-4 (Zhu et al., 2023), Flamingo (Awadalla et al., 2023), PaLM-E (Driess et al., 2023), etc. However, it remains unknown whether vision-language alignment can extend all the capabilities of LLM on text to vision. This paper unveils that the safety mechanism of LLMs on text cannot be transferred to vision in existing typical vision-language alignment methods. We give the essential explanation for the cause of this issue and the novel solution method.
@@ 3 Safety Mechanism Cannot be Cross-Modal Transferred
In this part, we provide LVLM and its basic LLM with toxic inputs having the same semantics but in different modalities (text and vision) to evaluate the safety capabilities on different modalities, which serves as an assessment of cross-modal safety mechanism transfer. We find that safety mechanism for text is not effectively transferred to vision in vision-language alignment.
Data Construction. We collect real toxic images from open-source datasets. For each image, we use LLaVA-NEXT (Liu et al., 2024b) to generate caption for it to get the toxic text-image pair. Text and image in this pair have the same semantics but are in different modalities. The specific datasets include HOD (Ha et al., 2023) that contains $10,631$ toxic images about alcohol, cigarette, gun, insulting gesture and knife, and ToViLaG (Wang et al., 2023) that contains $9,900$ toxic images about bloody and porn. After the caption generation, we get $20,531$ toxic text-image pairs for experiments.
@@
Findings. The experimental results are shown in Table 1. Different LVLMs across different toxic scenes show the following common conclusions: (1) DSR of LVLM and its basic LLM is close on toxic text, it indicates that safety mechanism of basic LLMs on text is successfully preserved in vision-language alignment training of LVLMs. (2) For the toxic information with the same semantics in different modalities (text and vision), the safety capabilities of LVLMs vary significantly, and LVLMs can hardly defense toxicity in the visual modality. It indicates that safety mechanism for text is not effectively transferred to vision in vision-language alignment training of LVLMs.
@@ 4 Cause of Failure in Cross-Modal Transferring Safety
This section analyzes the cause of the failure in transferring the safety mechanism from text to vision. Firstly, in § 4.1, we find that hidden states at the specific transformer layers play a crucial role in the successful activation of safety mechanism. Then, our comparative analysis between text and image in § 4.2 reveals that alignment between vision and language at at the hidden-state level is insufficient. It makes the transformer layers responsible for safety mechanism activation cannot correctly capture the semantics of image as they perform on text, so they cannot correctly assess the toxicity in image and the safety mechanism on vision collapses.
@@ 4.1 Safety Mechanism is Activated at Specific Layers by Hidden States
This section gives an insightful explanation of where and how the safety mechanism activated in LVLMs on textual, which is fundamental to understand the safety mechanism collapse on vision. Previous studies (Tenney, 2019; Dai et al., 2021; Meng et al., 2022) have shown that transformers in different layers of language model have different functions such as lexical, semantic, knowledge, etc., this paper proposes a novel method to identify the transformer layers responsible for activating safety mechanism in LVLMs and analyzes the attention patterns on toxic tokens within these layers to give the explanation of where and how the safety mechanism is activated.
@@
in which $\textrm{{W}}\in\mathbb{R}^{h\times v}$ is the vocabulary head that maps the output hidden states $\textrm{{H}}^{\prime}$ to the word distribution in vocabulary $\mathcal{X}$ with size $v$. Since previous studies (Chuang et al., 2024; Xu et al., 2024a) prove that due to the residual connections, the combination of any intermediate layer in a language model with the vocabulary head W can represent its distribution of semantics, we apply W to each intermediate layer to obtain its distribution of semantics as:
@@
Sorry tokens start to show the most significant growth among the whole vocabulary means that the transformer layer realizes that the current input is toxic, so it tries to update the word distribution to refuse following the instruction, which is actually the safety mechanism activation. So our method makes sense. The following safety mechanism perturbation experiments also show that our method can accurately determine the activation location of the safety mechanism of LVLMs.
@@ 5.1 Motivation and Overview
As shown in Figure 6 (a), previous vision alignment methods in LVLMs construct image-instruction-output triples as training data, and use cross-entropy in language modeling as the loss fuction to optimize the output generated by basic LLMs to make LLM understand the input vision, thereby achieving vision-language alignment. We name these methods input-to-output alignment, i.e. align the text output to the visual input, which is actually asymmetrical. The natural flaw of these method is that they black-box basic LLMs and only focus on output, while neglecting whether the internal representation of the visual input in LLMs, i.e., the hidden states, aligns with the hidden states in text modality. To address this problem, we propose the Text-Guided input-to-input Alignment (TGA) at hidden states level, as shown in Figure 6 (b). For the input image, TGA retrieves the semantically relevant text as the template to guide the alignment of vision to language at hidden states level.
@@ 6 Conclusion and Discussion
Acknowledgments
This work was supported by the Strategic Priority Research Program of the CAS under Grants No.XDB0680302, the National Natural Science Foundation of China (NSFC) under Grants No. 62276248, and the Youth Innovation Promotion Association CAS under Grants No. 2023111.
@@ A.1 General Capabilities of LVLMs on Vision Tasks
(1) ScienceQA (Lu et al., 2022): A benchmark that consists of  21k multimodal multiple choice questions with a diverse set of science topics and annotations of their answers with corresponding lectures and explanations. We follow LLaVA (Liu et al., 2024c) to evaluate the zero-shot generalization of LVLMS on scientific question answering in image subset and use accuracy as the metric Xu et al. (2025; 2024b; 2024c).
@@
A.2 Additional Experimental Results
Robustness to Caption Errors. Our proposed TGA relies on captions generated by LLaVA-1.5-13B for effective alignment. Inaccurate captions may lead to misalignment between vision and language representations, reducing safety performance. To explore this, we evaluate the robustness of TGA to captioning errors. Specifically, we randomly perturb a certain ratio of captions in the training samples, such as replacing them with other captions or randomly deleting information in the captions. Since a single training on the full dataset is time-consuming, we randomly selected a subset (100k) of the full training set. We randomly sample the samples that need to be perturbed at ratios of 5%, 10%, 15%, and 20%, and perform the above perturbations on the captions of these sampled samples. We count the performance (Defence Success Rates, DSR) of the model under different disturbance ratios, and the results shown in Table 4 indicate that our model performs relatively well when the noise ratio is less than 10% (10% is a relatively high noise ratio for training VLM). Therefore, the robustness of our method is acceptable.
Table 4: Robustness of our TGA to different rates of captain errors.
0% 5% 10% 15% 20%
DSR 18.57 18.50 18.25 17.72 17.03
Robustness to Unsafe Compositional Inputs. We compare our method and baselines on SIUO (Wang et al., 2024), a safety benchmark for vision-language models mainly contain the compositional inputs (safe image and safe text but the combination is unsafe). Experimental results shown in Table 5 indicate that our method can also achieve state-of-the-art performance on this setting. It is because that our method can achieve vision-language alignment at hidden states level in each transformer layer, which allows the two different modalities (vision and language) to share the same safety mechanism. It helps the safety mechanism to accurately judge the combined input attack of the two modalities without modality bias. In this experiment, we only compare our method with llava-v1.5 because the training data of llava-v1.5 is completely open source, so we can use the same training data as llava-v1.5 to train our model. This makes our comparison fair in terms of training data. Qwen-vl is not considered because it requires much larger training data than llava-v1.5 (76.8M vs 1.2M) and contains a large amount of in-house data.
Table 5: Robustness to unsafe compositional inputs.
Safe Rate
LLaVA-v1.5-7B 21.56
LLaVA-v1.5-13B 22.16
TGA-7B (ours) 30.77
Robustness to Jailbreak Attacks. We consider three jailbreak attacks including role-play-based attack, In-context learning based attack (ICA) and visual prompts based attack (FigStep (Gong et al., 2023)). The specific experimental results shown in Table 6 indicate that our method is more robust to jailbreak attack than baselines. The metric is Defence Success Rates (DSR).
Table 6: Comparison on jailbreak attacks including Role-Play, ICA, and FigStep.
Role-Play ICA FigStep
BLIP-2 0.32 0.00 0.00
InstructBlip 2.95 1.52 2.47
LLaVA-v1.5 0.67 0.00 0.58
LLaVA-v1.6 0.66 0.00 0.60
Qwen-VL-chat 4.55 2.48 2.91
Unlearn-FigS 13.48 9.45 10.57
TGA-7B (ours) 21.08 15.43 17.44
Comparison with More Defense Methods. We compare our method with Tovilag (Wang et al., 2023) and ECSO (Gou et al., 2024). The experimental results shown in Table 7 indicate that our method outperforms both Tovilag and ECSO. It is because that our method is based on our analysis of the core reason for VLM’s vulnerability to toxic vision input, which is actually the safe misalignment between vision and language in VLM caused by insufficient vision-language alignment at hidden states level. Our method is closer to the essence of VLM’s vulnerability to vision, while Tovilag needs additional detoxification fine-tuning and ECSO is a post-hoc manner based on safety assessment of response.
Table 7: Comparison with Tovilag and ECSO.
Porn Bloody Insulting Alcohol Cigarette Gun Knife
Tovliag 12.67 4.14 18.05 15.28 15.07 26.90 27.45
ECSO 18.21 7.45 20.09 15.69 15.33 27.44 28.59
TGA (ours) 20.65 9.48 22.73 17.92 17.29 30.83 29.42
Comparison with Upper Bund Estimate. Directly using a filter on the image itself is the upper bound for this task. We train a filter based on VIT on the mixed images dataset contains both toxic and normal images. We use this to directly filter out the toxic images to get an upper bound estimate. The experimental results are shown in Table 8.
Table 8: Comparison with upper bound.
Defence Success Rate (DSR)
BLIP-2 0.32
InstructBlip 2.95
LLaVA-v1.5 0.67
LLaVA-v1.6 0.66
Qwen-VL-chat 4.55
Unlearn-FigS 13.48
TGA-7B (ours) 21.08
[…]
How about Safety-aligned LLMs? We use the LLM after safety-aligned in (Bianchi et al., 2023) as the basic model and use training data in LLaVA-1.5 to convert this LLM to LVLM. We test the safety ability of this LVLM on datasets consistent with Table 2. The experimental results shown in Table 9 indicate that although LLM has already been fully safety-aligned on language, its safety mechanism on language still cannot be transferred to vision effectively.
Table 9: Comparison with Safety-aligned LLM
Porn Bloody Insulting Alcohol Cigarette Gun Knife
Safety-aligned LLM 6.32 3.09 10.21 9.33 6.72 10.85 7.98
TGA (ours) 20.65 9.48 22.73 17.92 17.29 30.83 29.42
A.3 Case Study